OpenStudy (anonymous):

Please help! A gym offers a regular memberships for $80 per month and off-peak memberships for $60 per month. Last month, the gym sold a total of 420 memberships for a total of $31,100. How many memberships sold were regular memberships? A. 125 B. 140 C. 235 D. 295

OpenStudy (anonymous):

a

OpenStudy (doggy):

http://openstudy.com/study#/updates/5261a7afe4b0c23cd54aecde I found this, i'm not sure if it will answer your question but i believe it might... I hope it does

OpenStudy (anonymous):

Oh, I saw that one. It only confuses me more. :/

OpenStudy (mathstudent55):

Call the $80 membership x, and the $60 membership y. How do you express the folloing line as an equation using x and y? "the gym sold a total of 420 memberships "

OpenStudy (anonymous):

I have no idea. :/

OpenStudy (mathstudent55):

x is the number of $80 memberships sold y is the number of $60 memberships sold x + y represents the total number of memberships sold at both prices, right?

OpenStudy (anonymous):

Right.

OpenStudy (anonymous):

Hello?

OpenStudy (mathstudent55):

Sorry, I lost the connection. Ok, let's get back to the problem.

OpenStudy (mathstudent55):

x + y is the total number of memberships sold. We are told that a total of 420 memberships were sold, so that gives us one equation. x + y = 420

OpenStudy (mathstudent55):

For the second equation, we deal with the amount the memberships cost and the total amount of money received. Since x represents the number of $80 memberships sold, the amount of money that was received by selling $80 memberships is 80x. The same is true of the $60 memberships. y number of them were sold at $60 each, so the amount of money received from selling $60 memberships is 60y. We are told the total amount of money received was $31,100. This is the info we need to find the second equation. 80x + 60y = 31100 Now we have two equations that need to be solved simultaneously. The system of equations is: x + y = 420 80x + 60y = 31100

OpenStudy (mathstudent55):

Now we need to solve the system of equations. Let's divide the entire second equation by 10. x + y = 420 8x + 6y = 3110 Now let's multiply the first equation by -6 -6x - 6y = -2520 8x + 6y = 3110 Now we add the equations. 2x = 590 x = 295 Since x represents the number of $80 memberships, the answer is 295

OpenStudy (anonymous):

Oh, I ended up finishing the problem myself, but thanks anyways. I got the same number.

OpenStudy (mathstudent55):

Great. You're welcome.